OpenStudy (anonymous):

Find the equation of the line that passes through the points (-2, 3) and (1, -6). Write your answer in Standard Form. Answer A: 3x + y = 3 B: 3x + y = -3 C: -3x – y = -3 D: -3x – y = 3

jimthompson5910 (jim_thompson5910):

Step 1 is to find the slope. How do we do this?

OpenStudy (anonymous):

y2-y1/x2-x1

jimthompson5910 (jim_thompson5910):

good, so what do you get for the slope?

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

y = mx + b y = -3x + b ... Plug in the slope you just found

jimthompson5910 (jim_thompson5910):

Now plug in either point. I'm going to pick (-2,3)

jimthompson5910 (jim_thompson5910):

y = -3x + b 3 = -3(-2) + b Solve for b

OpenStudy (anonymous):

b=-3

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

one sec, sry

jimthompson5910 (jim_thompson5910):

yes, b = -3 So the equation y = -3x + b becomes y = -3x - 3

OpenStudy (anonymous):

thankssssssssssssssssss

jimthompson5910 (jim_thompson5910):

Now convert to standard form